W(2W-5) = 180 need to know what W is.
Answer by AnlytcPhil(1806)

W(2W - 5) = 180 need to know what W is.

W(2W - 5) = 180 

Distribute to remove the parentheses:

            2W² - 5W = 180

Now get 0 on the right by adding -180 to both sides

      2W² - 5W - 180 = 0

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where x = W, a = 2; b = -5; c = -180

                      ________________ 
             -(-5) ± Ö(-5)²-4(2)(-180)
        W = ———————————————————————————
                     2(2) 
                   _______ 
              5 ± Ö25+1440
        W = ————————————————
                  4

                   ____ 
              5 ± Ö1465
        W = ————————————————
                  4

Using a calculator with the +, we get 10.8188296
Using a calculator with the -, we get -8.318829605

Edwin